OpenStudy (anonymous):

find the value of x for m(arc)AB=27 and m(arc)CD=49

OpenStudy (bradely):

incomplete question

OpenStudy (anonymous):

38

OpenStudy (bradely):

what is x

OpenStudy (anonymous):

38

OpenStudy (bradely):

figure was missed

OpenStudy (anonymous):

no 38 is the correct answer ,I just took a quiz and I got 76 from a different equation (x + x = measure of corresponding arcs, so 2x = measure of corresponding arcs) THIS IS WRONG FORMULA! http://openstudy.com/study#/updates/5016c6eae4b04dfc808af0bf <- ✓✓✓ this is correct formula below: x = 1/2*(measure of arc AB + measure of arc CD) x = 1/2*(27 + 49) x = 1/2*(76) x = 76/2 x = 38 ... which is choice A